# CAD-Unet: A Capsule Network-Enhanced Unet Architecture for Accurate Segmentation of COVID-19 Lung Infections from CT Images

### v1

@@ 3.4 Attention gate
The structure of the attention gate is elucidated in Fig.4. This module takes an input feature map $x\in R^{H\times W\times C_{x}}$ and a gate signal $g\in R^{H\times W\times C_{g}}$, where $x$ signifies the output feature map of the encoder $l$-th layer, and $g$ represents the output feature map of the decoder $(l+1)$-th layer. Subsequently, the feature map $x$ and the gate signal $g$ undergo linear transformations through $1\times 1$ convolutional layers, followed by batch normalization and element-wise summation. Following this, the ReLU activation function is applied, and another $1\times 1$ convolutional layer is utilized to derive spatial attention coefficients, denoted as Att for each pixel. Finally, these obtained attention coefficients are applied to the input feature map $x$, as expressed in equation (5):
@@
Fig. 4: Attention gate block, where $g$ is the gating signal and the $x$ is the input feature maps. $A$ is the obtained spatial attention, which is applied for all channels of the input feature maps $x$.
3.5 Dual decoder
To focus the network on the infected regions, we employ the dual decoder used in DAtt-Unet [3]. The dual decoder is designed to achieve simultaneous segmentation of both infections regions and lung regions. The objective is to guide the training process to explore the interior of the lung regions and differentiate between infected tissue and non-pulmonary tissue.
As shown in Fig.1, the decoder consists of two pathways: pathway 1 predicts the lung region, and pathway 2 predicts the infection region. The prediction results from pathway B serve as the final prediction target, while the existence of pathway A aims to guide the model focus on the lung region, as infection areas only occur within the lungs.
Specifically, except for the last layer, each layer in the decoder receives two inputs: the output from the previous layer of the decoder and the concatenated result of the attention gate output at the same layer. Similar to the encoder, these inputs undergo a ResBlock and an upsampling operation to become the current layer output. This output serves two purposes: 1) it is forwarded to the attention gate unit as the gating signal $g$, and 2) it is passed to the next layer of the decoder. The dual decoder process can be represented as equation (6):
@@ 3.6 Hybrid loss function
For CT lung lesion segmentation tasks, the complex morphology of the target region edges poses issue for accurate segmentation. Existing research has shown that edge information can provide valuable constraints to guide feature extraction for segmentation [28, 25, 27]. Therefore, we employ the hybrid loss function proposed in PDEAttUnet [3] to guide the network training, which consists of three components: 1) infection segmentation loss ($\mathcal{L}_{Inf}$) for the final task prediction, 2) lung segmentation loss ($\mathcal{L}_{Lung}$) to assist in the final prediction of lung regions, and 3) edge loss ($\mathcal{L}_{Edge}$) to enhance the segmentation of target edges. The hybrid loss function is formulated as equation (7):
@@
where $\alpha$, $\beta$, and $\gamma$ are the weights for the infection segmentation loss ($\mathcal{L}_{\text{Inf}}$), lung segmentation loss ($\mathcal{L}_{\text{Lung}}$), and edge loss ($\mathcal{L}_{\text{Edge}}$), respectively. $W$ and $H$ represent the width and height of the predicted mask, and $B$ is the batch size. The ground truth labels for infection and lung of pixel $i$ are denoted as $G_{R_{i}}\in\{0,1\}$ and $G_{L_{i}}\in\{0,1\}$, respectively. Additionally, $p_{i}$ and $s_{i}$ represent the prediction probabilities of infection and lung for pixel $i$, obtained from the decoders used for segmenting infection and lung, respectively. On the other hand, foredge loss function, $G_{E_{i}}\in\{0,1\}$ and $P_{E_{i}}$ are the ground truth label and the predicted probability of edge infection for pixel $i$, respectively. The ground-truth edge pixels are derived by applying morphological gradient to the ground-truth infection regions. Once the ground-truth edge pixels are obtained, the binary cross entropy (BCE) loss function is employed to compute the loss between the ground-truth edge pixels and their corresponding pixels in the mask map. The weight $\gamma$ of the edge loss is determined through experimental evaluation for each dataset.
4 Experiments and results
4.1 Dataset
@@
Dataset 4 CC-CCII segmentation dataset [26] 150 450
@@
The COVID-19 CT segmentation [17] dataset consists of 100 axial CT images (slices) from over 40 COVID-19 infected patients. It includes multi-class segmentation labels and will be used for the multi-class segmentation task.
Segmentation dataset nr.2 [17] contains 9 3D CT scans with a total of 829 slices. Among these slices, 373 have been annotated by radiologists with binary and multi-class segmentation labels for COVID-19 infection. This dataset will be used for both binary and multi-class segmentation tasks.
The COVID-19-CT-Seg dataset [12] comprises 20 COVID-19 CT scans, where all cases are confirmed COVID-19 infections. This dataset has been annotated by experienced radiologists and labeled for binary classification. Among the total of 3520 slices, 1844 slices are infected. It will be used for binary segmentation tasks.
The CC-CCII segmentation dataset [26] consists of 750 CT slices from 150 COVID-19 patients, manually annotated for background, lung field, GGO, and CON segmentation. It will be utilized for both binary and multi-class segmentation tasks.
@@ 4.2 Evaluation metrics
The model is evaluated by using the following metrics:
@@
    F1 score (F1-S): The F1-S is the harmonic mean of precision and recall, indicating the balance between the true positive rate and the false positive rate,
@@
    Recall(Rec): The Rec, also referred to as sensitivity or the true positive rate, quantitatively measures the proportion of true positive predictions among all actual positive instances,
@@
    Specificity (Spec): The Spec measures the proportion of true negative predictions among all actual negative instances,
@@
    Precision (Prec): The Prec represents the proportion of true positive predictions among all positive predictions made by the model,
@@ 4.3 Experimental setup
We conduct model training using the PyTorch framework. The experimental platform consists of an Intel Core i5 13600kf CPU, a single NVIDIA GeForce RTX 3090ti 24GB GPU, and the Windows 11 operating system. We utilize the Adam optimizer and set the batch size to 6, performing 120 epochs of training on the network. The initial learning rate is set to 1e-4 and is decayed by a factor of 0.1 after the 50th epoch and then again after the 90th epoch. Additionally, we apply data augmentation to the dataset, including random angle rotations ranging from -35 degrees to 35 degrees, as well as random horizontal and vertical flips. In terms of the hybrid loss function, the weights are set as 0.7, 0.3, and 1, respectively.
@@ 4.4 Experimental results
For the binary segmentation task, we compared our model with Unet, UNet++, Att-Unet, InfNet, SCOAT-Net, nCoVSegNet, and PDEAtt-Unet. Tables 2, 3, and 4 present the experimental results of the proposed model on datasets 2, 3, and 4, respectively. It can be observed that the proposed model achieved superior performance in most metrics across these datasets.
On dataset 2, our model outperformed the best-performing segmentation model, PDEAtt-Unet, with improvements of 4.23 in F1 score, 5.85 in IoU, and 4.90 in Recall. These significant improvements can be attributed to the parallel integration of the capsule network pathway and the U-Net encoder pathway in our model design. Capsule networks excel at capturing and preserving hierarchical relationships between objects, while U-Net is adept at extracting local features from images. Through the pathway coupling mechanism, our model effectively combines the advantages of both networks, resulting in more accurate segmentation of complex COVID-19 lung lesions.
On dataset 3, our model achieved improvements of 1.02 in F1 score, 1.33 in IoU, and 2.40 in Recall compared to PDEAtt-Unet. These enhancements further demonstrate the effectiveness of our model design, particularly in handling the diversity and complexity of different datasets. By integrating the capsule network and U-Net encoder in parallel, our model can more effectively identify and segment various lung lesions caused by COVID-19.
On dataset 4, our model also showed improvements in F1 score, IoU, and Recall compared to PDEAtt-Unet, with increases of 0.1, 0.15, and 1.94, respectively. Although these improvements are relatively smaller, they still indicate the consistency and stability of our model across different datasets. This consistency is due to the pathway coupling mechanism in our model, which allows efficient information flow between the two network pathways, enhancing segmentation accuracy and robustness.
Model F1-S IoU Rec Spec Prec
InfNet 40.86 25.68 35.11 99.06 48.88
SCOATNet 69.39 53.13 56.67 99.83 89.46
nCoVSegNet 55.41 38.32 41.10 99.81 85.01
Unet 66.38 49.68 52.15 99.87 91.30
UNet++ 59.29 42.13 43.23 99.93 94.32
Att-Unet 59.32 42.17 45.62 99.79 84.79
PDEAtt-Unet 77.60 63.40 80.69 99.30 74.74
Ours 81.61 68.93 82.79 99.55 80.46
@@
Model F1-S IoU Rec Spec Prec
InfNet 72.91 57.37 66.48 99.78 80.72
SCOATNet 71.02 55.06 61.35 99.84 84.29
nCoVSegNet 62.48 45.44 50.16 99.86 82.85
Unet 65.82 49.05 53.32 99.88 85.97
UNet++ 69.89 53.72 60.69 99.82 82.39
Att-Unet 68.67 52.29 58.34 99.84 83.45
PDEAtt-Unet 75.52 60.67 81.94 99.52 70.03
Ours 78.13 64.11 86.01 99.6 71.56
@@
Model F1-S IoU Rec Spec Prec
InfNet 72.19 56.49 63.62 99.86 83.44
SCOATNet 77.57 63.25 70.66 99.86 85.12
nCoVSegNet 78.06 64.02 74.69 99.82 81.74
Unet 75.49 60.63 69.34 99.84 82.84
UNet++ 77.45 63.20 71.55 99.85 84.42
Att-Unet 77.27 62.96 70.48 99.87 85.51
PDEAtt-Unet 79.38 65.81 82.52 99.72 76.47
Ours 80.90 67.93 87.19 99.66 75.46
@@
In the multi-class segmentation task, we merge dataset 1 and dataset 2 into one multi-class segmentation dataset. On this combined dataset, we compare the proposed model with Unet, Att-Unet, Unet++, CopleNet, AnamNet, SCOAT-Net, and D-TrAttUnet among other models. Table 5 presents the experimental results of the proposed model for multi-class segmentation. It can be observed that the model achieved the best segmentation performance, outperforming the previous state-of-the-art segmentation model D-TrAttUnet by 2.79 in F1 score and 3.25 in IoU for the GGO classification. For the CON classification, the model surpasses the previous best segmentation model D-TrAttUnet by 2.00 in F1 score and 1.94 in IoU. Additionally, unlike D-TrAttUnet, the model does not use multiple layers of transformers, resulting in significantly reducing the number of model parameters by almost an order of magnitude compared to D-TrAttUnet.
Model GGO Consolidation Params
F1-S IoU F1-S IoU
Unet 65.81 49.06 31.35 19.26 $\sim$
Att-Unet 64.81 47.97 39.04 24.48 $\sim$
Unet++ 65.69 48.92 31.31 18.75 $\sim$
CopleNet 60.44 43.33 29.70 17.90 $\sim$
AnamNet 65.10 48.36 31.97 19.18 $\sim$
SCOATNET 65.77 49.09 43.52 27.83 $\sim$
D-TrAttUnet 67.25 50.67 55.24 38.16 105.81M
Ours 70.71 54.69 59.61 42.47 15.05M

### v2

@@ 3.4 Attention gate
The structure of the attention gate is elucidated in Fig.4. This module processes an input feature map $x\in R^{H\times W\times C_{x}}$ and a gate signal $g\in R^{H\times W\times C_{g}}$, where $x$ signifies the output feature map of the encoder $l$-th layer, and $g$ represents the output feature map of the decoder $(l+1)$-th layer. Subsequently, the feature map $x$ and the gate signal $g$ undergo linear transformations through $1\times 1$ convolutional layers, followed by batch normalization and element-wise summation. Following this, the ReLU activation function is applied, and another $1\times 1$ convolutional layer is utilized to derive spatial attention coefficients, denoted as $Att$ for each pixel. Finally, these obtained attention coefficients are applied to the input feature map $x$, as expressed in equation (5):
@@
Fig. 4: Attention gate block, where $g$ is the gating signal and $x$ is the input feature maps. $A$ is the spatial attention obtained, which is applied to all channels of the input feature maps $x$.
3.5 Dual decoders
To focus the network on the infected regions, we employ the dual decoders used in DAtt-Unet [3]. The dual decoders is designed to achieve simultaneous segmentation of both infections regions and lung regions. The objective is to guide the training process to explore the interior of the lung regions and differentiate between infected tissue and non-pulmonary tissue.
As shown in Fig.1, the decoder consists of two pathways: pathway 1 predicts the lung region, and pathway 2 predicts the infection region. The prediction results from pathway 2 serve as the final prediction target, while the existence of pathway 1 aims to guide the model to focus on the lung region, as infection areas only occur within the lungs.
Specifically, except for the last layer, each layer in the decoder receives two inputs: the output from the previous layer of the decoder and the concatenated result of the attention gate output at the same layer. Similar to the encoder, these inputs undergo a ResBlock and an upsampling operation to produce the current layer’s output. This output serves two purposes: 1) it is forwarded to the attention gate unit as the gating signal $g$, and 2) it is passed to the next layer of the decoder. The dual decoders process can be represented as equation (6):
@@ 3.6 Hybrid loss function
For CT lung lesion segmentation tasks, the complex morphology of the target region edges poses issue for accurate segmentation. Existing works have shown that edge information can provide valuable constraints to guide feature extraction for segmentation [30, 27, 29]. Therefore, we employ the hybrid loss function proposed in PDEAttUnet [3] to guide the network training, which consists of three components: 1) infection segmentation loss ($\mathcal{L}_{Inf}$) for the final task prediction, 2) lung segmentation loss ($\mathcal{L}_{Lung}$) to assist in the final prediction of lung regions, and 3) edge loss ($\mathcal{L}_{Edge}$) to enhance the segmentation of target edges. The hybrid loss function is formulated as equation (7):
@@
where $\alpha$, $\beta$, and $\gamma$ are the weights for the infection segmentation loss ($\mathcal{L}_{\text{Inf}}$), lung segmentation loss ($\mathcal{L}_{\text{Lung}}$), and edge loss ($\mathcal{L}_{\text{Edge}}$), respectively. $W$ and $H$ represent the width and height of the predicted mask, and $B$ is the batch size. The ground truth labels for infection and lung of pixel $i$ are denoted as $G_{R_{i}}\in\{0,1\}$ and $G_{L_{i}}\in\{0,1\}$, respectively. Additionally, $p_{i}$ and $s_{i}$ represent the prediction probabilities of infection and lung for pixel $i$, obtained from the decoders used for segmenting infection and lung, respectively. On the other hand, foredge loss function, $G_{E_{i}}\in\{0,1\}$ and $P_{E_{i}}$ are the ground truth label and the predicted probability of edge infection for pixel $i$, respectively. The ground-truth edge pixels are derived by applying morphological gradient to the ground-truth infection regions. Once the ground-truth edge pixels are obtained, the binary cross entropy (BCE) loss function is employed to compute the loss between the ground-truth edge pixels and their corresponding pixels in the mask map. The specific values of the weights $\alpha$, $\beta$, and $\gamma$ will be described in the experimental setup section.
@@
Dataset 4 CC-CCII segmentation dataset [28] 150 750
@@
Name Total slices Non-infected Mild Intermediate Severe
Dataset 1 100 1 52 27 20
Dataset 2 829 457 280 48 44
Dataset 3 1844 0 1582 193 69
Dataset 4 750 201 462 72 15
Table 2: Distribution of CT slices by infection severity levels across the datasets.
4 Experiments and results
4.1 Dataset
@@
The COVID-19 CT Segmentation dataset [19] comprises 100 axial CT slices collected from over 40 patients diagnosed with COVID-19. This dataset provides multi-class segmentation labels and is intended for use in multi-class segmentation tasks.
Segmentation dataset nr.2 [19] includes 9 3D CT scans, comprising a total of 829 slices. Of these, 373 slices have been annotated by radiologists with both binary and multi-class segmentation labels for COVID-19 infection. This dataset will be utilized for binary and multi-class segmentation tasks.
The COVID-19-CT-Seg dataset [14] includes 20 CT scans from confirmed COVID-19 cases. It covers a wide range of infection percentages (from 0.01% to 59%). A total of 1844 CT slices are annotated with binary infection labels, a meticulous process carried out by a multi-tiered team ranging from junior annotators to senior radiologists. This dataset is intended for binary segmentation tasks.
The CC-CCII segmentation dataset [28] includes 750 CT slices from 150 COVID-19 patients, with manual annotations for background, lung field, GGO, and CON segmentation. This dataset will be employed for both binary and multi-class segmentation tasks.
Furthermore,to evaluate the quality and potential biases of the datasets used in this study, we provide a detailed breakdown of the distribution of infection severity levels across the CT slices in each dataset. Table 2 summarizes the number of CT slices categorized by different infection severity levels, as defined by [28].: mild, defined as less than three ground-glass opacity (GGO) lesions of size less than 3 cm; intermediate, defined as a lesion area exceeding 25% of the entire lung field; and severe, defined as a lesion area exceeding 50% of the entire lung field. As observed in the table, non-infected slices in dataset 2 account for more than half of the data (55%). To mitigate the impact of this bias on model training and validation, we performed random undersampling to reduce their proportion to approximately 30%. This adjustment helps prevent the model from being skewed towards the majority class (non-infected slices) and improves its ability to learn meaningful features from infected cases. It is important to note that the random undersampling process was carefully implemented to preserve the overall diversity and representativeness of the dataset.
@@ 4.2 Evaluation metrics
The model is evaluated by using the following metrics.
@@
    F1 score (F1-S): The F1 score is the harmonic mean of precision and recall, indicating the balance between the true positive rate and the false positive rate,
@@
    Recall(Rec): The Recall, also referred to as sensitivity or the true positive rate, quantitatively measures the proportion of true positive predictions among all actual positive instances,
@@
    Specificity (Spec): The Specificity measures the proportion of true negative predictions among all actual negative instances,
@@
    Precision (Prec): The Precision represents the proportion of true positive predictions among all positive predictions made by the model,
@@ 4.3 Experimental setup
We conduct model training using Python 3.7.12 and the PyTorch 1.13.0 framework. The experimental platform consists of an Intel Core i5 13600KF CPU, a single NVIDIA GeForce RTX 3090 Ti 24GB GPU, and the Windows 11 operating system. The libraries used include NumPy 1.21.6, OpenCV 4.7.0, MONAI 1.1.0, among others. The Adam optimizer is adopted for training, with a batch size of 6 and a total of 120 epochs. The initial learning rate is set to 1e-4 and is decayed by a factor of 0.1 after the 50th epoch and again after the 90th epoch. Additionally, we apply data augmentation to the dataset, including random angle rotations ranging from -35 degrees to 35 degrees, as well as random horizontal and vertical flips. In terms of the hybrid loss function, the weights $\alpha$, $\beta$, and $\gamma$ are set to 0.7, 0.3, and 1, respectively. The main experimental settings are summarized in Table 3.
Category Specification
CPU Intel Core i5 13600KF
GPU
NVIDIA GeForce RTX 3090 Ti
24GB
Operating System Windows 11
RAM 32GB DDR4
Programming Language Python 3.7.12
Framework PyTorch 1.13.0
CUDA Version 11.7
cuDNN Version 8.5.0
Additional Libraries
NumPy 1.21.6, OpenCV 4.7.0,
MONAI 1.1.0
Epoch 120
Batchsize 6
Optimizer Adam
Learning rate 1E-04
Table 3: Summary of experimental setup.
@@ 4.4 Experimental results
In this study, we conduct comprehensive experiments to evaluate the proposed model for both binary and multi-class segmentation tasks of COVID-19 lung CT lesions. All the parameters of the compared methods are set according to the recommendations in the original references or adjusted to achieve the best visual and quantitative results. In addition, nnUNet is widely recognized as a robust baseline for medical image segmentation. To further assess the effectiveness of our approach, we incorporate the proposed architecture into nnUNet, which we term nnCAD-UNet, and compare its performance with the original nnUNet on both segmentation tasks. All experimental results are reported as the mean ± standard deviation over five independent runs. The following sections provide a qualitative analysis of the two tasks.
4.4.1 Binary segmentation task
For the binary segmentation task, we compare the proposed model with U-Net, U-Net++, Att-UNet, InfNet, SCOAT-Net, nCoVSegNet, PDEAtt-UNet, and D-TrAttUnet. To evaluate the impact of the hybrid loss on the segmentation performance of different models, we conducted experiments by incorporating the hybrid loss into U-Net, U-Net++, Att-UNet, SCOAT-Net, and nCoVSegNet (note that InfNet, PDEAtt-UNet, and D-TrAttUnet use their own specific loss functions). The experimental results on datasets 2, 3, and 4 are presented in Tables 4, 5, and 6.
On dataset 2, the proposed model outperforms all compared models across multiple metrics, including nnUNet and nnCAD-UNet. Specifically, compared to the best-performing dedicated COVID-19 segmentation model, nCoVSegNet with hybrid loss, the proposed model improves the F1 score, IoU, and Recall by 4.25, 3.65, and 5.9, respectively. When compared to nnUNet, the overall best-performing model among all methods, the proposed model achieves improvements of 1.87, 0.54, and 10.56 in these three metrics. However, integrating the proposed model into nnUNet leads to performance degradation across all three metrics, likely due to the specific infection characteristics in dataset 2 and the relatively large proportion of non-infection CT slices, which may not align well with nnUNet’s preprocessing strategy.
Compared to dataset 2, dataset 3 exhibits overall improved segmentation performance, with F1 scores surpassing 70% for most models. The general improvement in IoU and Recall, along with lower standard deviations, suggests that the lesions in dataset 3 are more distinct, making segmentation relatively less challenging. On this dataset, the proposed model continues to demonstrate strong segmentation performance. Specifically, compared to the best-performing model among the compared methods, AttUNet, the proposed model achieves improvements of 1.33, 1.81, and 10.12 in F1 score, IoU, and Recall, respectively. Additionally, nnUNet, with its automatically configured architecture, continues to exhibit strong general segmentation capabilities, outperforming all non-nnUNet-based models. However, integrating the proposed model into nnUNet results in improvements of 1.15 and 1.64 in F1 score and IoU, respectively, compared to the original nnUNet, while the Recall metric decreases by 2.51. This indicates that the integration may enhance certain aspects of segmentation accuracy but could compromise the model’s ability to fully capture true positive cases.
Among the three datasets, dataset 4 yields the highest segmentation performance across all models. The F1 scores range from 71.61% (U-Net) to 85.56% (nnCAD-UNet), with IoU values also being the highest among the datasets. Specifically, the proposed model achieves the best results, outperforming SCOATNet, the top-performing model among non-nnUNet-based architectures, by 1.15, 1.6, and 9.4 in terms of F1 score, IoU, and Recall respectively. After integrating this architecture into nnUNet, compared to the original nnUNet, improvements are observed in all three metrics. This indicates that, whether in its original form or when integrated into nnUNet, our proposed Capsule Network-Enhanced Unet Architecture exhibits excellent adaptability and robustness.
In addition, analysis of the experimental results comparing models with and without hybrid loss across all datasets reveals that while the F1 score remains relatively unchanged or experiences a slight decrease, the Recall on most models improves. This aligns with the priorities in medical diagnostics, where minimizing false negatives is of utmost importance. A higher Recall ensures the identification of more true positive cases, thus reducing the risk of missed diagnoses. Although the F1 score does not show a substantial increase, the enhanced Recall improves the model’s ability to detect lesions, which is critical for early COVID-19 detection and diagnosis. Therefore, we conclude that the hybrid loss function provides a valuable improvement for the binary segmentation task of COVID-19 lesions.
Model F1 Score IoU Recall Spec Prec
Unet 45.23±13.00 30.15±11.05 37.88±12.63 99.39±0.45 60.47±16.54
UNet++ 52.50±9.96 36.23±9.42 43.54±11.96 99.55±0.23 69.67±9.96
Att-Unet 63.63±8.74 46.77±6.02 48.96±7.37 99.89±0.13 91.41±7.22
InfNet 47.96±5.30 31.69±4.40 42.92±4.78 99.17±0.38 55.54±11.26
SCOATNet 62.06±7.32 45.40±7.85 47.93±8.70 99.77±0.12 83.94±6.96
nCoVSegNet 75.35 ±3.21 60.55±4.12 68.16±5.74 99.72±0.08 84.73±2.97
PDEAtt-Unet 68.74±8.55 46.27±9.00 64.84±12.10 99.09±0.24 61.57±6.87
D-TrAttUnet 74.44±2.38 59.34±3.13 67.97±4.96 99.68±0.07 61.74±2.51
Unet(Hybrid Loss) 41.82±13.56 27.44±11.71 30.75±12.01 99.69±0.2 69.89±16.77
UNet++(Hybrid Loss) 51.37±1.72 34.58±1.57 57.72±14.52 98.52±1.01 51.74±14.16
Att-Unet(Hybrid Loss) 50.05±9.74 33.94±8.65 38.04±9.12 99.72±0.11 74.64±10.76
SCOATNet(Hybrid Loss) 54.71±18.95 24.23±8.66 48.57±23.58 99.57±0.30 70.63±10.88
nCoVSegNet(Hybrid Loss) 75.57±2.73 60.81±3.56 72.83±4.05 99.55±0.12 78.77±4.44
CAD-Unet 79.82±1.89 64.46±2.64 78.73±3.81 99.58±0.10 81.18±3.70
nnUnet 77.95±2.60 63.92±3.47 68.17±3.25 99.85±0.02 91.06±1.45
nnCAD-Unet 77.15±1.07 62.81±1.41 64.43±1.58 99.88±0.01 96.17±0.35
@@ 4.4.1 Binary segmentation task
Model F1 Score IoU Recall Spec Prec
Unet 74.38±2.77 57.22±4.46 66.59±7.70 99.80±0.03 79.74±5.14
UNet++ 75.99±1.18 62.08±3.39 73.95±3.29 99.77±0.02 79.37±2.40
Att-Unet 76.64±0.92 62.09±0.99 72.86±1.55 99.84±0.07 83.09±3.16
InfNet 74.67±0.66 59.58±0.84 80.02±1.54 99.59±0.04 70.04±1.90
SCOATNet 75.33±1.93 60.47±2.46 73.35±4.16 99.75±0.06 77.72±3.60
nCoVSegNet 72.92±1.34 57.40±1.66 74.94±3.95 99.62±0.06 70.34±2.62
PDEAtt-Unet 74.32±0.87 59.15±1.12 84.82±1.22 99.49±0.03 66.16±1.34
D-TrAttUnet 71.15±2.00 55.26±2.41 81.74±2.29 99.44±0.08 63.08±3.24
Unet(Hybrid Loss) 74.35±2.71 59.24±3.38 68.19±4.36 99.82±0.02 81.93±1.70
UNet++(Hybrid Loss) 75.36±0.70 60.47±0.90 83.58±1.33 99.56±0.02 68.62±1.04
Att-Unet(Hybrid Loss) 76.36±1.30 62.04±1.77 72.58±2.80 99.80±0.02 81.10±1.57
SCOATNet(Hybrid Loss) 71.72±1.28 55.93±1.56 75.38±4.55 99.59±0.09 68.76±3.91
nCoVSegNet(Hybrid Loss) 72.97±1.07 57.46±1.34 75.89±3.59 99.62±0.06 70.48±2.77
CAD-Unet 77.97±0.72 63.90±0.98 82.98±1.38 99.65±0.04 73.58±2.25
nnUnet 82.86±0.21 70.81±0.29 84.67±0.65 99.85±0.01 81.48±0.61
nnCAD-Unet 84.01±0.18 72.45±0.27 82.16±0.47 99.90±0.00 86.03±0.46
@@
Model F1 Score IoU Sens Spec Prec
Unet 71.61±1.43 55.79±1.73 63.36±1.52 99.83±0.03 82.41±3.31
UNet++ 78.94±0.71 65.22±0.96 75.94±1.17 99.80±0.02 82.22±1.75
Att-Unet 77.20±0.57 64.95±0.80 77.18±1.01 99.81±0.14 83.49±0.99
InfNet 70.92±0.53 54.95±0.64 71.53±1.33 99.63±0.02 70.36±1.37
SCOATNet 79.68±0.32 66.23±0.45 76.77±1.41 99.80±0.02 82.85±1.13
nCoVSegNet 76.60±0.32 62.08±0.43 80.07±0.87 99.63±0.01 72.61±0.97
PDEAtt-Unet 77.76±0.48 62.29±0.63 83.89±3.03 99.58±0.06 70.91±2.88
D-TrAttUnet 77.16±0.41 62.82±0.54 86.12±2.73 99.55±0.04 69.97±1.49
Unet(Hybrid Loss) 76.76±1.41 62.31±1.86 70.96±2.88 99.83±0.01 83.71±1.05
UNet++(Hybrid Loss) 77.38±0.57 63.11±0.75 84.68±1.28 99.59±0.03 71.27±1.57
Att-Unet(Hybrid Loss) 77.55±0.56 64.44±0.78 77.97±0.97 99.78±0.09 82.37±1.98
SCOATNet(Hybrid Loss) 78.92±0.46 65.18±0.64 82.85±1.30 99.67±0.03 75.39±1.85
nCoVSegNet(Hybrid Loss) 76.78±0.29 62.32±0.39 80.79±0.93 99.64±0.02 73.18±1.11
CAD-Unet 80.83±0.20 67.83±0.28 86.17±1.72 99.69±0.02 76.71±1.06
nnUnet 85.16±0.17 74.16±0.26 81.85±0.18 99.88±0.01 88.75±0.49
nnCAD-Unet 85.56±0.20 74.77±0.30 82.58±0.56 99.88±0.01 88.78±0.73
@@
4.4.2 Multi-class segmentation task
For the multi-class segmentation task, we merge dataset 1 and dataset 2 into a single multi-class segmentation dataset. On this combined dataset, we compare the proposed model with U-Net, Att-UNet, U-Net++, CopleNet, AnamNet, SCOAT-Net, D-TrAttUnet, and nnUNet. The experimental results, as shown in Table 6, demonstrate that the proposed model also performs well on this task. Specifically, compared to D-TrAttUnet, the best-performing non-nnUNet-based architecture, the proposed model achieves a slight improvement of 0.1 in the F1 score for the GGO class and a more significant improvement of 1.67 for the more challenging CON class. After integrating the proposed model into nnUNet, although the F1 score for the GGO class decreases by 0.26, it achieves a substantial improvement of 2.26 for the CON class, along with a lower standard deviation, indicating greater model stability.
The experimental results collectively demonstrate that the proposed model achieves competitive performance in both binary and multi-class segmentation tasks. This is attributed to the capsule network’s ability to identify complex lesion postures and the ingenious pathway coupling mechanism. The capsule network effectively captures the unique geometric and spatial features of lesions, while the pathway coupling mechanism facilitates efficient information fusion. Their synergy leads to accurate lesion detection, as evidenced by the experimental results across four datasets in both tasks. These elements are crucial for enhancing the segmentation of COVID-19 lung CT lesions, offering a reliable approach in medical image analysis.
Model GGO CON
F1 score IoU F1 score IoU
Unet 49.75±8.52 33.54±7.58 43.23±4.98 27.70±3.91
Att-Unet 69.06±0.88 53.46±1.04 54.61±1.62 37.58±1.53
Unet++ 65.69±1.29 48.92±14.2 31.31±6.67 18.75±4.73
CopleNet 60.44±1.54 43.33±1.61 29.70±10.29 17.90±7.52
AnamNet 65.10±3.56 48.36±3.82 31.97±6.12 19.18±4.36
SCOATNET 65.77±3.28 49.09±3.56 43.52±1.67 27.83±1.38
D-TrAttUnet 70.61±1.01 54.58±1.21 57.94±2.30 40.82±3.66
CAD-Unet 70.71±0.44 54.69±0.52 59.61±0.74 42.47±0.75
nnUnet 73.39±1.02 57.97±1.26 60.30±2.20 43.19±2.23
nnCAD-Unet 73.13±0.22 57.65±0.27 62.56±0.85 45.53±0.90